Prove by using distance formula that the points P(1,2,3) , Q(-1,-1,-1) and R(3,5,7) are collinear.

Grade:11

1 Answers

Gosula Madhukar
7 years ago
let P=(1,2,3) and Q=(-1,-1,-1) and R=(3,5,7).then by distance formula pq+pr=qr.
here pq=29^1/2
qr=116^1/2
pr=29^1/2
29^1/2+29^1/2=116^1/2
there fore given points are collinear.